OpenStudy (shortypants):

Let f​(x​)equals=x^2+4x and g​(x​)equals=1−x. Find ​f(2)minus−​g(2​). ​f(2​)minus−​g(2​)equals=___ ​(Simplify your​ answer.)

OpenStudy (shortypants):

would this be 2x^2+4(2)-1-(2)

OpenStudy (shortypants):

i mean (2^2)+4(2)-1-(2)

OpenStudy (shortypants):

if so is the answer 9

OpenStudy (mathstudent55):

"=" means "equals" "-" means "minus" There is no need to write =equals or -minus which just confuses.

OpenStudy (shortypants):

sorry that was just a copy paste error from the program im using

OpenStudy (brydenwright):

If you put 2 into f(x), you get 12, or 4+8. When you plug in 2 into g(x), you get 1-2, or -1. 12--1 is the same as 12+1, so the answer would be 13.

OpenStudy (shortypants):

oh i see you solved both sides on there own before subtracting them?

OpenStudy (brydenwright):

Yeah, because it's really (x^2+4x)-(1-x)

OpenStudy (shortypants):

ok thank you.

OpenStudy (mathstudent55):

\(f​(x​)=x^2+4x\) and \(g​(x​)=1−x\) There are different ways of finding your answer. Here is one way. First, find an expression for f(x) - g(x). Then let x = 2 and evaluate. \(f​(x​) - g(x) =x^2+4x - (1−x)\) \(f​(x​) - g(x) =x^2+4x - 1 + x\) \(f​(x​) - g(x) =x^2+5x - 1)\) \(f​(2​) - g(2) =2^2+5(2) - 1 = 4 + 10 - 1 = 13\) Here is another way. First evaluate f(2) and g(2) then do the subtraction. \(f​(x​)=x^2+4x\) \(f​(2​)=2^2+4(2) =4 + 8 = 12\) \(g​(x​)=1−x\) \(g​(2​)=1−2 = -1\) \(f(2) - g(2) = 12 - (-1) = 12 + 1 = 13\)